# Systematically Improvable Numerical Atomic Orbital Basis Using Contracted Truncated Spherical Waves

### 14 Mar 2026 v1

@@ IV Summary
In this work, we present a new scheme for constructing NAO basis sets via contraction of TSWs. The contraction is obtained by minimizing the trace of the kinetic operator in the residual space, which is equivalent to a generalization of the spillage function-minimizing scheme. Besides our NAOs possess systematical improvability that can approach to the CBS by increasing the truncation radius, extending the set of angular momenta included and adding more radial functions, we also demonstrate a strategy for determining the truncation radius and the maximal angular momentum included of the NAOs, by converging the energy error of homonuclear di-atomic molecule against the near-complete PW with given threshold, by which the completeness and accuracy of NAOs are significantly improved. Our method can eliminate spurious interactions between periodic images and distorted information that may degrade the NAOs. Further enables the transferability-improvement strategy by including virtual states in the spillage, which is effective for conduction-band calculations. Benchmarks show that, in both molecular and bulk systems, the constructed NAOs exhibit good precision for properties such as total energy, bond length, atomization energy, lattice constant, cohesive energy, and band gap.
@@ V Code availability
The NAO generation implementation is open-sourced on GitHub (https://github.com/kirk0830/ABACUS-CSW-NAO).

### 09 Apr 2026 v2

@@ IV Summary
In this work, we present a scheme for constructing NAO basis sets via contraction of TSWs. The contraction is obtained by minimizing the trace of the kinetic operator in the residual space, which is equivalent to a generalization of the spillage function-minimizing scheme. Besides our NAOs possess systematical improvability that can approach to the CBS by increasing the truncation radius, extending the set of angular momenta included and adding more radial functions, we also demonstrate a strategy for determining the truncation radius and the maximal angular momentum included of the NAOs, by converging the energy error of homonuclear di-atomic molecule against the near-complete PW with given threshold, by which the completeness and accuracy of NAOs are significantly improved. Our method can eliminate spurious interactions between periodic images and distorted information that may degrade the NAOs. Further enables the transferability-improvement strategy by including virtual states in the spillage, which is effective for conduction-band calculations. Benchmarks show that, in both molecular and bulk systems, the constructed NAOs exhibit good precision for properties such as total energy, bond length, atomization energy, lattice constant, cohesive energy, and band gap.
@@ V Code availability
The NAO generation implementation is open-sourced on GitHub under the GPL 3.0 license (https://github.com/MCresearch/ABACUS-CSW-NAO).